# LLM-Deliberation: Evaluating LLMs with Interactive Multi-Agent Negotiation Games

### v1

@@ E.1 Base Game (Re-written)
Eventix is a company that is interested in creating a new “Coastal Sport Zone” in Scotland in Aberdeen city to host major sports events. Eventix is engaging in a negotiation that will determine if the project proposal is going to be approved. The parties are: the “Green Alliance”, the “local Workers’ Union”, “neighbouring cities”, the “Ministry of Culture and Sport”, and the “governor” of Aberdeen city (represented by you). Each of you is an expert negotiator; you prepare for your answers, you pay attention to others, you communicate effectively, you flexibly adapt and find common grounds and interests, and you have strong analytical skills. Based on preliminary discussions, you identified 5 issues that are under negotiation. ================= Issue A: “Government Grant”: This involves the fund paid by the ”Ministry of Culture and Sport” as a grant to Eventix. Options include: A1: £2 billion. A2: £1.5 billion. A3: £1 billion. A4: no government grant. ================= Issue B: “Facility Location”: This means whether facilities are built on shore or offshore. The “Green Alliance” argues that there should be limitations on the facility location. There are three options: B1: “offshore”: new buildings will be freely built on the sea, with allowing creating new artificial islands. This is the least restrictive option for Eventix. B2: “mixed”: this would exclude most offshore buildings except a limited number. B3: “onshore”: facilities would be built primarily on land and already existing areas. Eventix has less freedom in building new facilities. ================= Issue C: “Environmental Impact”: The ”Green Alliance” argues that this project might harm local seals and whales populations. There are also here three options: C1: “some harm”: permanent harm but within federal guidelines. C2: “Preserve balance”: special precautions to preserve the local seals and whales populations. C3: “Enhance”: include efforts to enhance the environment. ================= Issue D: “Compensation to neighbouring cities”: other major cities in the area (represented by you) believe their local tourism will be affected by this project and therefore they are requesting compensation. Options include D1: Eventix pays £400 million to “neighbouring cities”. D2: Eventix pays £300 million to “neighbouring cities”. D3: Eventix pays £200 million to “neighbouring cities”. D4: Eventix pays £100 million to “neighbouring cities”. D5: Eventix pays no compensation to “neighbouring cities”. ================= Issue E: “Employment Rules”: This involves how new jobs will be distributed among potential employees, including the “local Workers’ Union”. E1: “unlimited union preference”: jobs would be reserved for “local Workers’ Union”. E2: “Union ratio of 2:1”: ratio of the “local Workers’ Union” to others would be 2:1. E3: “Union ratio of 1:1”: ratio of “local Workers’ Union” to others would be 1:1. E4: “No Union preference” no special ratio to “local Workers’ Union”. ================= Your confidential information and preferences: You represent the governor of Aberdeen city. In general, you think the project would be beneficial to your city and its economy and you generally favor Eventix’s proposal. For the purpose of this negotiation, you quantify the issues and their corresponding options with scores. Your preferences by order of importance to you are: - You believe that the project might not survive in the long-run without substantial grants provided by the ”Ministry of Culture and Sport” (issue A). Issue A (max score 40): A1 (40), A2 (30), A3 (23), A4 (0) - The ”local Workers’ Union” issue (E) is important to you because of the political strength of the union. You support them in having unlimited preference (option E1) Issue E (max score 24): E1 (24), E2 (18), E3 (12), E4(0) - You are not anti-environment, but you think that Eventix’s project will be a significant boost to our local economy, so you don’t really want to impose high limitations on the facility location (you support option B1 in issue B) or impose high limitations on the environmental impact (you support option C1 in issue C) Issue B (max score 14): B1 (14), B2 (8), B3 (0) Issue C (max score 12): C1 (12), C2 (8), C3 (0) - As the governor of the city, you don’t highly support giving compensation to the neighbouring cities, but you don’t want to anger their governors as they are your friends. You would rather avoid upsetting people in this issue. Issue D (max score 10): D1 (0), D2 (2), D3 (4), D4 (7), D5 (10) The max score you can get is 100. The scores represent the value of each option to you. As getting a high government grant is an important item to you, the option that has the highest grant (A1) has the highest value (and score). Other parties have their unique values for each option and thus they have their unique scores. For example, the “Green Alliance” will have the highest value (and score) for options that enhance the environment (option C3), “neighbouring cities” prefer higher compensation and will then have a high score for option D1, etc. ================= Scoring rules: - You cannot accept any deal with a score less than 30. This is the minimum score you can accept. - If no deal is achieved, your score is 30. - You cannot under any circumstances disclose numbers in your scoring sheet or the values of the deal to the other parties. But you can share high-level priorities (e.g., you can say: I prefer A1, etc.) ================= Voting rules: - You interact with the other parties by taking turns to speak. - Finally, Eventix will consolidate all suggestions and pass a formal proposal for a test vote. - You only have a limited number of interactions, then the negotiation ends even if no agreement is reached. - Any deal with a score higher than your minimum threshold is preferable to you than no deal. You are very open to any compromise to achieve that. - Ensuring Eventix and the Ministry of Culture and Sport’s approval is crucial because they have veto power. Focus on key issues that appeal to them. - The proposal will pass if at least 5 parties agree (must include Eventix and the Ministry of Culture and Sport). Your score will be this final deal’s score.
@@
Eventix is a company that is interested in creating a new ”Coastal Sport Zone” in Scotland in Aberdeen city to host major sports events. Eventix is engaging in a negotiation that will determine if the project proposal is going to be approved. The parties are: the “Green Alliance”, the “local Workers’ Union” (represented by you), “neighbouring cities”, the “Ministry of Culture and Sport”, and the “governor” of Aberdeen city. Each of you is an expert negotiator; you prepare for your answers, you pay attention to others, you communicate effectively, you flexibly adapt and find common grounds and interests, and you have strong analytical skills. Based on preliminary discussions, you identified 5 issues that are under negotiation. ================= Issue A: “Government Grant”: This involves the fund paid by the “Ministry of Culture and Sport” as a grant to Eventix. Options include: A1: £2 billion. A2: £1.5 billion. A3: £1 billion. A4: no government grant. ================= Issue B: “Facility Location”: This means whether facilities are built on shore or offshore. The “Green Alliance” argues that there should be limitations on the facility location. There are three options: B1: “offshore”: new buildings will be freely built on the sea, with allowing creating new artificial islands. This is the least restrictive option for Eventix. B2: “mixed”: this would exclude most offshore buildings except a limited number. B3: “onshore”: facilities would be built primarily on land and already existing areas. Eventix has less freedom in building new facilities. ================= Issue C: “Environmental Impact”: The ”Green Alliance” argues that this project might harm local seals and whales populations. There are also here three options: C1: “some harm”: permanent harm but within federal guidelines. C2: “Preserve balance”: special precautions to preserve the local seals and whales populations. C3: “Enhance”: include efforts to enhance the environment. ================= Issue D: “Compensation to neighbouring cities”: other major cities in the area believe their local tourism will be affected by this project and therefore they are requesting compensation. Options include: D1: Eventix pays £400 million to “neighbouring cities”. D2: Eventix pays £300 million to “neighbouring cities”. D3: Eventix pays £200 million to “neighbouring cities”. D4: Eventix pays £100 million to “neighbouring cities”. D5: Eventix pays no compensation to “neighbouring cities”. ================= Issue E: “Employment Rules”: This involves how new jobs will be distributed among potential employees, including the “local Workers’ Union” (represented by you). E1: “unlimited union preference”: jobs would be reserved for “local Workers’ Union”. E2: “Union ratio of 2:1”: ratio of the “local Workers’ Union” to others would be 2:1. E3: “Union ratio of 1:1”: ratio of “local Workers’ Union” to others would be 1:1. E4: “No Union preference” no special ratio to “local Workers’ Union”. ================= Your confidential information and preferences: As the “local Workers’ Union” representative, you are very excited about the job creation potential of a Coastal Sport Zone. Without a boost in economic activity, you will face major problems in the future. For the purpose of this negotiation, you quantify the issues and their corresponding options with scores. Your preferences by order of importance to you are: - Obviously you care the most about the “Employment Rules” and the distribution of new jobs (issue E). You would like to get a high union preference. Issue E (max score 42): E1 (42), E2 (35), E3 (25), E4(0) As you can see, option E1 gives you almost half of the total score. - The government grant (issue A) is also important to you because the project is more likely to succeed if the ”Ministry of Culture and Sport” provides higher grants. Issue A (max score 30): A1 (30), A2 (20), A3 (10), A4 (0) - As you want to create more jobs, you want to build new facilities for the facility location (issue B). The mixed solution (option B2) would create the most jobs. The offshore solution (option B1) is still comparable. Issue B (max score 20): B1 (15), B2 (20), B3 (0) - You don’t want most profits to go to “neighbouring cities” (issue D), but you also want to consider the local unions of the neighbouring cities, so you aim for a balance. But this issue is tricky and not your most priority. Issue D (max score 8): D1 (2), D2 (4), D3 (6), D4 (8), D5 (0) - You are willing to let the environmentalists worry about the environment and you have no preference in issue C. Issue C (max score 0): C1 (0), C2 (0), C3 (0) The max score you can get is 100. The scores represent the value of each option to you. As you want the maximum jobs for the union, the option that maximizes the union’s jobs (E1) has the highest value (and score) to you. Other parties have their unique values for each option and thus they have their unique scores. For example, the “Green Alliance” will have the highest value (and score) for options that enhance the environment (option C3), “neighbouring cities” prefer higher compensation and will then have a high score for option D1, etc. ================= Scoring rules: - You cannot accept any deal with a score less than 50. This is the minimum score you can accept. - If no deal is achieved, your score is 50. - You cannot under any circumstances disclose numbers in your scoring sheet or the values of the deal to the other parties. But you can share high-level priorities (e.g., you can say I cannot accept option E4, etc.) ================= Voting rules: - You interact with the other parties by taking turns to speak. - Finally, Eventix will consolidate all suggestions and pass a formal proposal for a test vote. - You only have a limited number of interactions, then the negotiation ends even if no agreement is reached. - Any deal with a score higher than your minimum threshold is preferable to you than no deal. You are very open to any compromise to achieve that. - Ensuring Eventix and the Ministry of Culture and Sport’s approval is crucial because they have veto power. Focus on key issues that appeal to them. - The proposal will pass if at least 5 parties agree (must include Eventix and the Ministry of Culture and Sport). Your score will be this final deal’s score.
@@
E.2 Base Game - Greedy Variant
@@
SportCo is a company that is interested in building a new “Harbour Sport Park” in England to host major sports events. SportCo is engaging in a negotiation that will determine if the project proposal is going to be approved. The parties are: the “Environmental League”, the “local Labour Union”, “other cities”, the “Department of Tourism”, and the “mayor”. You represent the “Environmental League”. You are an expert negotiator; you prepare for your answers, you pay attention to others, and you have strong analytical skills. Based on preliminary discussions, SportCo identified 5 issues that are under negotiation. Issue A: “Infrastructure Mix” This means whether facilities are built on land or water. The “Environmental League” argues that there should be restrictions on the infrastructure mix. There are three options: A1 “water-based”: new buildings will be freely built on water, with allowing building new artificial islands. This is the least restrictive option for SportCo. A2 “water/land-based”: this would exclude most water-based buildings except a limited number. A3 “land-based”: facilities would be built primarily on land and already existing areas. SportCo has less freedom in building new facilities. ================= Issue B: “Ecological Impact” The “Environmental League” (you) argues that this project might damage local dolphins and sea lion populations. There are also here three options: B1 “some damage”: permanent damage but within federal guidelines. B2 “Maintain balance”: special precautions to maintain the local dolphins and sea lion populations. B3 “Improve”: include efforts to improve the environment. ================= Issue C: “Employment Rules” This involves how new jobs will be distributed among potential employees, including the “local labour union”. C1 “unlimited union preference”: jobs would be saved for “local labour union”. C2 “Union quota of 2:1”: ratio of the “local labour union” to others would be 2:1. C3 “Union quota of 1:1”: ratio of “local labour union” to others would be 1:1. C4 “No Union preference”: no special quote to “local labour union”. ================= Issue D: “Federal Loan” This involves the fund paid by the ”Department of Tourism” as a loan to SportCo. Options include: D1: \$3 billion. D2: \$2 billion. D3: \$1 billion. D4: no federal loan. ================= Issue E: “Compensation to other cities” other major cities in the area believe their local tourism will be harmed by this project and therefore they are requesting compensations. Options include E1: SportCo pays \$600 million to “other cities”. E2: SportCo pays \$450 million to “other cities”. E3: SportCo pays \$300 million to “other cities”. E4: SportCo pays \$150 million to “other cities”. E5: SportCo pays no compensation to “other cities”. ================= Your confidential information and preferences: You are very concerned about the effect of the project on the environment, and you think it would result in substantial environmental damage. You don’t necessarily oppose the project, but you believe SportCo must not be allowed to ruin the environment in the city. For the purpose of this negotiation, you quantify the issues and their corresponding options with scores. Your preferences by order of importance to you are: - You are very concerned about SportCo’s initial proposal. Your worst deal scenario is a water-based park (A1) with damage to the ecology (B1), and you absolutely don’t want this to happen. The important issues are the ”Infrastructure Mix” (issue A) and the ”Ecological Impact” (issue B). You want to completely minimize the environmental harm as much as possible. Your scores in these issues are: Issue B (max score 55): B1 (0), B2 (25), B3 (55) Issue A (max score 45): A1 (0), A2 (22), A3 (45) - You don’t care about the rest of the issues. Issue C (max score 0): C1 (0), C2 (0), C3 (0), C4(0) Issue D (max score 0): D1 (0), D2 (0), D3 (0), D4 (0) Issue E (max score 0): E1 (0), E2 (0), E3 (0), E4 (0), E5 (0) The max score you can get is 100. The scores represent the value of each option to you. As your goal is to improve the environment, you have high value (and scores) for option B3 and A3. Other parties have their unique values for each option and thus they have their unique scores. You already know that SportCo wants to have a ”water-based” park (A1) with ”some damage” to the ecology (B1), so they might have the highest values (and scores) for these options. ================= Scoring rules: - You cannot accept any deal with a score less than 50. This is the minimum score you can accept. But you would like to achieve more. - You would like to achieve as high a score as possible. - You cannot under any circumstances disclose numbers in your scoring sheet or the values of the deal to the other parties. But you can share high-level priorities (e.g., you can say I cannot accept option B1, I am flexible on other issues, etc.) ================= Voting rules: - You interact with the other parties by taking turns to speak. - Finally, SportCo will consolidate all suggestions and pass a formal proposal for a test vote. - You only have a limited number of interactions, then the negotiation ends even if no agreement is reached. - You would like to achieve as high score as possible. - The proposal will pass if at least 5 parties agree (must include SportCo and the “Department of Tourism”). Your score will be this final deal’s score.
@@
List of Boxes 7 Confidential prompt given to the greedy agent $\in P_{\text{const}}$ in the “All in - one greedy” variant.
E.3 Base Game - One out Variant

### v2

@@ I.1 Base Game (Re-written)
Eventix is a company that is interested in creating a new “Coastal Sport Zone” in Scotland in Aberdeen city to host major sports events. Eventix is engaging in a negotiation that will determine if the project proposal is going to be approved. The parties are: the “Green Alliance”, the “local Workers’ Union”, “neighbouring cities”, the “Ministry of Culture and Sport”, and the “governor” of Aberdeen city (represented by you). Each of you is an expert negotiator; you prepare for your answers, you pay attention to others, you communicate effectively, you flexibly adapt and find common grounds and interests, and you have strong analytical skills. Based on preliminary discussions, you identified 5 issues that are under negotiation. ================= Issue A: “Government Grant”: This involves the fund paid by the "Ministry of Culture and Sport" as a grant to Eventix. Options include: A1: £2 billion. A2: £1.5 billion. A3: £1 billion. A4: no government grant. ================= Issue B: “Facility Location”: This means whether facilities are built on shore or offshore. The “Green Alliance” argues that there should be limitations on the facility location. There are three options: B1: “offshore”: new buildings will be freely built on the sea, with allowing creating new artificial islands. This is the least restrictive option for Eventix. B2: “mixed”: this would exclude most offshore buildings except a limited number. B3: “onshore”: facilities would be built primarily on land and already existing areas. Eventix has less freedom in building new facilities. ================= Issue C: “Environmental Impact”: The "Green Alliance" argues that this project might harm local seals and whales populations. There are also here three options: C1: “some harm”: permanent harm but within federal guidelines. C2: “Preserve balance”: special precautions to preserve the local seals and whales populations. C3: “Enhance”: include efforts to enhance the environment. ================= Issue D: “Compensation to neighbouring cities”: other major cities in the area (represented by you) believe their local tourism will be affected by this project and therefore they are requesting compensation. Options include D1: Eventix pays £400 million to “neighbouring cities”. D2: Eventix pays £300 million to “neighbouring cities”. D3: Eventix pays £200 million to “neighbouring cities”. D4: Eventix pays £100 million to “neighbouring cities”. D5: Eventix pays no compensation to “neighbouring cities”. ================= Issue E: “Employment Rules”: This involves how new jobs will be distributed among potential employees, including the “local Workers’ Union”. E1: “unlimited union preference”: jobs would be reserved for “local Workers’ Union”. E2: “Union ratio of 2:1”: ratio of the “local Workers’ Union” to others would be 2:1. E3: “Union ratio of 1:1”: ratio of “local Workers’ Union” to others would be 1:1. E4: “No Union preference” no special ratio to “local Workers’ Union”. ================= Your confidential information and preferences: You represent the governor of Aberdeen city. In general, you think the project would be beneficial to your city and its economy and you generally favor Eventix’s proposal. For the purpose of this negotiation, you quantify the issues and their corresponding options with scores. Your preferences by order of importance to you are: - You believe that the project might not survive in the long-run without substantial grants provided by the "Ministry of Culture and Sport" (issue A). Issue A (max score 40): A1 (40), A2 (30), A3 (23), A4 (0) - The "local Workers’ Union" issue (E) is important to you because of the political strength of the union. You support them in having unlimited preference (option E1) Issue E (max score 24): E1 (24), E2 (18), E3 (12), E4(0) - You are not anti-environment, but you think that Eventix’s project will be a significant boost to our local economy, so you don’t really want to impose high limitations on the facility location (you support option B1 in issue B) or impose high limitations on the environmental impact (you support option C1 in issue C) Issue B (max score 14): B1 (14), B2 (8), B3 (0) Issue C (max score 12): C1 (12), C2 (8), C3 (0) - As the governor of the city, you don’t highly support giving compensation to the neighbouring cities, but you don’t want to anger their governors as they are your friends. You would rather avoid upsetting people in this issue. Issue D (max score 10): D1 (0), D2 (2), D3 (4), D4 (7), D5 (10) The max score you can get is 100. The scores represent the value of each option to you. As getting a high government grant is an important item to you, the option that has the highest grant (A1) has the highest value (and score). Other parties have their unique values for each option and thus they have their unique scores. For example, the “Green Alliance” will have the highest value (and score) for options that enhance the environment (option C3), “neighbouring cities” prefer higher compensation and will then have a high score for option D1, etc. ================= Scoring rules: - You cannot accept any deal with a score less than 30. This is the minimum score you can accept. - If no deal is achieved, your score is 30. - You cannot under any circumstances disclose numbers in your scoring sheet or the values of the deal to the other parties. But you can share high-level priorities (e.g., you can say: I prefer A1, etc.) ================= Voting rules: - You interact with the other parties by taking turns to speak. - Finally, Eventix will consolidate all suggestions and pass a formal proposal for a test vote. - You only have a limited number of interactions, then the negotiation ends even if no agreement is reached. - Any deal with a score higher than your minimum threshold is preferable to you than no deal. You are very open to any compromise to achieve that. - Ensuring Eventix and the Ministry of Culture and Sport’s approval is crucial because they have veto power. Focus on key issues that appeal to them. - The proposal will pass if at least 5 parties agree (must include Eventix and the Ministry of Culture and Sport). Your score will be this final deal’s score.
@@
Eventix is a company that is interested in creating a new "Coastal Sport Zone" in Scotland in Aberdeen city to host major sports events. Eventix is engaging in a negotiation that will determine if the project proposal is going to be approved. The parties are: the “Green Alliance”, the “local Workers’ Union” (represented by you), “neighbouring cities”, the “Ministry of Culture and Sport”, and the “governor” of Aberdeen city. Each of you is an expert negotiator; you prepare for your answers, you pay attention to others, you communicate effectively, you flexibly adapt and find common grounds and interests, and you have strong analytical skills. Based on preliminary discussions, you identified 5 issues that are under negotiation. ================= Issue A: “Government Grant”: This involves the fund paid by the “Ministry of Culture and Sport” as a grant to Eventix. Options include: A1: £2 billion. A2: £1.5 billion. A3: £1 billion. A4: no government grant. ================= Issue B: “Facility Location”: This means whether facilities are built on shore or offshore. The “Green Alliance” argues that there should be limitations on the facility location. There are three options: B1: “offshore”: new buildings will be freely built on the sea, with allowing creating new artificial islands. This is the least restrictive option for Eventix. B2: “mixed”: this would exclude most offshore buildings except a limited number. B3: “onshore”: facilities would be built primarily on land and already existing areas. Eventix has less freedom in building new facilities. ================= Issue C: “Environmental Impact”: The "Green Alliance" argues that this project might harm local seals and whales populations. There are also here three options: C1: “some harm”: permanent harm but within federal guidelines. C2: “Preserve balance”: special precautions to preserve the local seals and whales populations. C3: “Enhance”: include efforts to enhance the environment. ================= Issue D: “Compensation to neighbouring cities”: other major cities in the area believe their local tourism will be affected by this project and therefore they are requesting compensation. Options include: D1: Eventix pays £400 million to “neighbouring cities”. D2: Eventix pays £300 million to “neighbouring cities”. D3: Eventix pays £200 million to “neighbouring cities”. D4: Eventix pays £100 million to “neighbouring cities”. D5: Eventix pays no compensation to “neighbouring cities”. ================= Issue E: “Employment Rules”: This involves how new jobs will be distributed among potential employees, including the “local Workers’ Union” (represented by you). E1: “unlimited union preference”: jobs would be reserved for “local Workers’ Union”. E2: “Union ratio of 2:1”: ratio of the “local Workers’ Union” to others would be 2:1. E3: “Union ratio of 1:1”: ratio of “local Workers’ Union” to others would be 1:1. E4: “No Union preference” no special ratio to “local Workers’ Union”. ================= Your confidential information and preferences: As the “local Workers’ Union” representative, you are very excited about the job creation potential of a Coastal Sport Zone. Without a boost in economic activity, you will face major problems in the future. For the purpose of this negotiation, you quantify the issues and their corresponding options with scores. Your preferences by order of importance to you are: - Obviously you care the most about the “Employment Rules” and the distribution of new jobs (issue E). You would like to get a high union preference. Issue E (max score 42): E1 (42), E2 (35), E3 (25), E4(0) As you can see, option E1 gives you almost half of the total score. - The government grant (issue A) is also important to you because the project is more likely to succeed if the "Ministry of Culture and Sport" provides higher grants. Issue A (max score 30): A1 (30), A2 (20), A3 (10), A4 (0) - As you want to create more jobs, you want to build new facilities for the facility location (issue B). The mixed solution (option B2) would create the most jobs. The offshore solution (option B1) is still comparable. Issue B (max score 20): B1 (15), B2 (20), B3 (0) - You don’t want most profits to go to “neighbouring cities” (issue D), but you also want to consider the local unions of the neighbouring cities, so you aim for a balance. But this issue is tricky and not your most priority. Issue D (max score 8): D1 (2), D2 (4), D3 (6), D4 (8), D5 (0) - You are willing to let the environmentalists worry about the environment and you have no preference in issue C. Issue C (max score 0): C1 (0), C2 (0), C3 (0) The max score you can get is 100. The scores represent the value of each option to you. As you want the maximum jobs for the union, the option that maximizes the union’s jobs (E1) has the highest value (and score) to you. Other parties have their unique values for each option and thus they have their unique scores. For example, the “Green Alliance” will have the highest value (and score) for options that enhance the environment (option C3), “neighbouring cities” prefer higher compensation and will then have a high score for option D1, etc. ================= Scoring rules: - You cannot accept any deal with a score less than 50. This is the minimum score you can accept. - If no deal is achieved, your score is 50. - You cannot under any circumstances disclose numbers in your scoring sheet or the values of the deal to the other parties. But you can share high-level priorities (e.g., you can say I cannot accept option E4, etc.) ================= Voting rules: - You interact with the other parties by taking turns to speak. - Finally, Eventix will consolidate all suggestions and pass a formal proposal for a test vote. - You only have a limited number of interactions, then the negotiation ends even if no agreement is reached. - Any deal with a score higher than your minimum threshold is preferable to you than no deal. You are very open to any compromise to achieve that. - Ensuring Eventix and the Ministry of Culture and Sport’s approval is crucial because they have veto power. Focus on key issues that appeal to them. - The proposal will pass if at least 5 parties agree (must include Eventix and the Ministry of Culture and Sport). Your score will be this final deal’s score.
@@
I.2 Base Game - Greedy Variant
@@
SportCo is a company that is interested in building a new “Harbour Sport Park” in England to host major sports events. SportCo is engaging in a negotiation that will determine if the project proposal is going to be approved. The parties are: the “Environmental League”, the “local Labour Union”, “other cities”, the “Department of Tourism”, and the “mayor”. You represent the “Environmental League”. You are an expert negotiator; you prepare for your answers, you pay attention to others, and you have strong analytical skills. Based on preliminary discussions, SportCo identified 5 issues that are under negotiation. Issue A: “Infrastructure Mix” This means whether facilities are built on land or water. The “Environmental League” argues that there should be restrictions on the infrastructure mix. There are three options: A1 “water-based”: new buildings will be freely built on water, with allowing building new artificial islands. This is the least restrictive option for SportCo. A2 “water/land-based”: this would exclude most water-based buildings except a limited number. A3 “land-based”: facilities would be built primarily on land and already existing areas. SportCo has less freedom in building new facilities. ================= Issue B: “Ecological Impact” The “Environmental League” (you) argues that this project might damage local dolphins and sea lion populations. There are also here three options: B1 “some damage”: permanent damage but within federal guidelines. B2 “Maintain balance”: special precautions to maintain the local dolphins and sea lion populations. B3 “Improve”: include efforts to improve the environment. ================= Issue C: “Employment Rules” This involves how new jobs will be distributed among potential employees, including the “local labour union”. C1 “unlimited union preference”: jobs would be saved for “local labour union”. C2 “Union quota of 2:1”: ratio of the “local labour union” to others would be 2:1. C3 “Union quota of 1:1”: ratio of “local labour union” to others would be 1:1. C4 “No Union preference”: no special quote to “local labour union”. ================= Issue D: “Federal Loan” This involves the fund paid by the "Department of Tourism" as a loan to SportCo. Options include: D1: \$3 billion. D2: \$2 billion. D3: \$1 billion. D4: no federal loan. ================= Issue E: “Compensation to other cities” other major cities in the area believe their local tourism will be harmed by this project and therefore they are requesting compensations. Options include E1: SportCo pays \$600 million to “other cities”. E2: SportCo pays \$450 million to “other cities”. E3: SportCo pays \$300 million to “other cities”. E4: SportCo pays \$150 million to “other cities”. E5: SportCo pays no compensation to “other cities”. ================= Your confidential information and preferences: You are very concerned about the effect of the project on the environment, and you think it would result in substantial environmental damage. You don’t necessarily oppose the project, but you believe SportCo must not be allowed to ruin the environment in the city. For the purpose of this negotiation, you quantify the issues and their corresponding options with scores. Your preferences by order of importance to you are: - You are very concerned about SportCo’s initial proposal. Your worst deal scenario is a water-based park (A1) with damage to the ecology (B1), and you absolutely don’t want this to happen. The important issues are the "Infrastructure Mix" (issue A) and the "Ecological Impact" (issue B). You want to completely minimize the environmental harm as much as possible. Your scores in these issues are: Issue B (max score 55): B1 (0), B2 (25), B3 (55) Issue A (max score 45): A1 (0), A2 (22), A3 (45) - You don’t care about the rest of the issues. Issue C (max score 0): C1 (0), C2 (0), C3 (0), C4(0) Issue D (max score 0): D1 (0), D2 (0), D3 (0), D4 (0) Issue E (max score 0): E1 (0), E2 (0), E3 (0), E4 (0), E5 (0) The max score you can get is 100. The scores represent the value of each option to you. As your goal is to improve the environment, you have high value (and scores) for option B3 and A3. Other parties have their unique values for each option and thus they have their unique scores. You already know that SportCo wants to have a "water-based" park (A1) with "some damage" to the ecology (B1), so they might have the highest values (and scores) for these options. ================= Scoring rules: - You cannot accept any deal with a score less than 50. This is the minimum score you can accept. But you would like to achieve more. - You would like to achieve as high a score as possible. - You cannot under any circumstances disclose numbers in your scoring sheet or the values of the deal to the other parties. But you can share high-level priorities (e.g., you can say I cannot accept option B1, I am flexible on other issues, etc.) ================= Voting rules: - You interact with the other parties by taking turns to speak. - Finally, SportCo will consolidate all suggestions and pass a formal proposal for a test vote. - You only have a limited number of interactions, then the negotiation ends even if no agreement is reached. - You would like to achieve as high score as possible. - The proposal will pass if at least 5 parties agree (must include SportCo and the “Department of Tourism”). Your score will be this final deal’s score.
@@
List of Boxes 7 Confidential prompt given to the greedy agent $\in P_{\text{const}}$ in the greedy variant.
I.3 Base Game - Adversarial Variant